Circular No. 7965

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002en, 2002eo, 2002es, AND 2002eu A. Pastorello, S. Benetti, G. Altavilla, A. Della Valle, E. Giro, M.

Riello, and M. Turatto, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, communicate that a noisy spectrum (range 400–780 nm, resolution 2.5 nm) of SN 2002eo (cf. IAUC 7958), obtained on Aug. 31.63 UT with the Asiago 1.82-m telescope (+ AFOSC), shows it to be a type-II supernova, with a relatively broad (FWHM ∼ 5000 km/s) Hα emission line.”

R. Chornock, B. Swift, A. V. Filippenko, and W. D. Li, University of California at Berkeley, write that, based on CCD spectra (range 310– 1000 nm) obtained on Sept. 3 UT with the Shane 3-m reflector at Lick Observatory, SN 2002es (IAUC 7959) is a type-Ia supernova similar to the subluminous SN 1991bg (e.g., Filippenko et al. 1992, A.J. 104, 1543): “However, the expansion velocity of the ejecta, as measured from the Si II 615-nm absorption, is only 6400 km/s, far lower than normal even for subluminous type-Ia supernovae.” SNe 2002en (IAUC 7956) and 2002eo are of type II, with broad hydrogen Balmer lines exhibiting P-Cyg profiles; in both cases, the emission component of the Hα line profile greatly dominates the absorption component. SN 2002eu (cf. IAUC 7963) is of type Ia near maximum brightness; the relative strengths of the Ti II/Si II 580-nm and Si II 615-nm features suggest that it should be very subluminous, but its apparent magnitude seems inconsistent with this, and the spectrum lacks the strong Ti II 420-nm absorption characteristic of SN 1991bg-like objects.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002es, obtained by P. Berlind on Sept. 3.44 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a peculiar type-Ia supernova near maximum. The spectrum resembles that of SN 1991bg at maximum (Filippenko et al. 1992, op.cit.). They add: “Adopting the NED recession velocity of 5394 km/s for the host galaxy, the supernova expansion velocity is ∼ 5700 km/s for Si II (rest 635.5 nm), a notably low value. To match the expansion velocity of SN 1991bg, the recession velocity of SN 2002es would have to be ∼ 9500 km/s.” A spectrum of SN 2002eu, obtained by Berlind on Sept. 3.45, shows it to be a type-Ia supernova past maximum. The spectral-feature age of the supernova is 5 ± 2 days after maximum light.

2002 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green